# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

### v4

@@ I Introduction
Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [9, 10, 11]. Early methods, such as the Linear Inverted Pendulum Model (LIPM) [12], allowed for simplified representations of the dynamics involved in bipedal movement, enabling easier analysis and control. As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion. While model-based approaches offer rapid convergence and predictive capabilities, they often fall short in dynamically complex and uncertain environments, where adaptability is key.
@@
To address these questions, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical. These frameworks enable robots to generate diverse gaits, including walking, running, and turning, while also allowing them to navigate complex terrains, such as stairs, slopes, and deformable surfaces.
@@
Current progress shows that we are still far from achieving a unified framework, despite two mainstream types of framework structure. Although bipedal locomotion tasks have evolved significantly in the past decade. Earlier frameworks primarily focused on basic walking stabilization and simple terrain adaptation, whereas recent developments have introduced complex tasks such as dynamic stair climbing, parkour-inspired locomotion, and loco-manipulation. Notably, the DARPA Robotics Challenge (DRC) has established a structured benchmark for bipedal humanoid locomotion and manipulation [39], emphasizing real-world deployment and robustness.
@@
The survey is organised as follows: Section II focuses on end-to-end frameworks, categorised by learning approaches. Section III details hierarchical frameworks, classified into three main types. Section IV presents the expected unified frameworks, which build upon proposed approaches while highlighting the challenges associated with their implementation. Section V addresses potential future research directions. Finally, Section VI concludes the survey.
Figure 2: Classification of DRL-based control schemes.
@@ II End-to-end framework
TABLE I: Summary and comparison of reference-based and reference-free Learning approaches for end-to-end framework. The dashed line in the implementation flow chat refers to optional.
Methods Works Capabilities Characteristic Advantages and Disadvantages Implementation Flow Chart
Residual learning [42] [43] Forward walk undirectional walk Add residual term to the known motor positions at the current time step. A: Fast convergence speed D: Require high-quality predefined reference, limit to specific motions, and lack robustness to complicated terrains. Refer to caption
Guided learning [44] [41] [45] [31] Forward walk Versatile walk Versatile jump Versatile motions Mimic the predefined reference and directly specifies joint-level command. A: Accelerate the learning process and robust to the terrains. D: Limited to the predefined motions and lack adaptability to unforeseen changes in environment. Refer to caption
Reference-free learning [28] [46] [29] Periodic motions Stepping stones walk Visual walk Learn locomotion skills from scratch without any prior knowledge. A: High potential for gait exploration, high robust to the complicated terrain. D: Requires intensive reward shaping for gait patterns and relatively expensive computational training resources. Refer to caption
00footnotetext: Forward Walk involves the bipeds walking straight ahead. Unidirectional Walk enables the bipeds to move both forward and backward within a range of desired velocities. Omni-Walk grants the bipeds the ability to walk in any direction. Versatile Walk allows the bipeds to walk forward, backward, turn, and move sideways, providing extensive movement capabilities. Periodic Motions entails the execution of various repeated gait patterns, such as walking, hopping, or galloping. Versatile Jump refers to jump towards different desired targets. Versatile Motions cover performing a broad array of motions, both periodic and aperiodic such as jumping.
@@ III Hierarchy framework
Control Scheme Works Module Characteristic Advantages and Disadvantages
Deep Planning Hybrid Scheme [33] [53] [54] Deep planning + ID Deep planning + ID-QP Deep planning + WPG HL policy is learned to guide the LL controller to complete locomotion and navigation tasks. A: Enhanced command tracking capabilities, generalised across different platforms, sampling efficiency, and robust. D: Complicate system and communication between layers, require precise model, lack generalisation regarding different tasks.
Feedback DRL Control Hybrid Scheme [55] [35] [36] Gait library + feedback policy Footstep planner + feedback policy Model-based planner + feedback policy LL feedback policy receives non-learned HL planner as input to achieve locomotion skills. A: Short inference times, robust, navigation locomotion capabilities, interpretability. D: Complicated system and communication between layers, reducing sampling efficiency.
Learned Hierarchy Framework [21] [38] [56] HL policy + LL policy HL policy + LL policy HL policy + LL policy Both HL planner and LL feedback controller are learned. LL policy focuses on basic locomotion skills; on the other side, HL policy learn navigation skills. A: provides layer flexibility, where each layer can be independently retrained and reused; alleviates the challenges associated with training an end-to-end policy. D: inefficienty sim-to-real, complicated interface between layers, training expensively.
@@ III-A Deep planning hybrid scheme
Figure 3: Hierarchy Control Scheme Diagram: presents a hierarchical control framework for a bipedal robot, consisting of a basic scheme and three variations. (1) Basic Scheme: The framework starts with a Task Command, followed by a HL Planner, and a LL Controller, which ultimately drives the Robot. Each module can be replaced by a learned policy, introducing adaptability in different control layers. (2) Variations (from left to right): (a) A deep planning hybrid scheme, where the HL planner is learned. (b) A Feedback DRL control hybrid scheme, with a learned LL controller. (c) A Learned hierarchy control scheme, where both layers are learned.
@@ III-C Learned hierarchy framework
IV Towards a Unified Framework
Previous Sections II and III outlined two distinct paradigms for controlling bipedal robot locomotion: end-to-end and hierarchical frameworks. While both paradigms have demonstrated significant success in addressing specific locomotion tasks, their ability to generalise across diverse robot morphologies, environmental conditions, and locomotion objectives remains limited. Consequently, this limitation motivates us to revisit the central question posed in Section I:
Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks? If not, how should such a framework be designed, and what features or components should it include?
To address this, we first explicitly define a unified framework for bipedal locomotion as a control architecture capable of supporting consistent learning, transfer, and execution across arbitrary robot morphologies, terrains, and task definitions, ultimately aiming to develop genuinely generalist locomotion policies that achieve robust zero-shot or few-shot generalisation to previously unseen scenarios, embodiments, and objectives.
Evaluating existing approaches against this definition reveals that neither the end-to-end nor hierarchical control paradigms currently meet these unified criteria. Typically, these methods require significant redesign or retraining when introduced to new morphologies, tasks, or environmental conditions, highlighting the practical gap towards achieving genuine unification.
To explore pathways towards unified frameworks, we identify two currently foreseeable and potentially viable directions that build upon the structural foundations of the previously discussed paradigms—end-to-end and hierarchical. These two directions offer distinct but complementary strategies for advancing towards scalable, generalisable locomotion architectures.
  • Bipedal Foundation Models (BFMs) – large-scale, pre-trained models that map perception directly to action through representation learning. Trained on diverse data across tasks and embodiments, BFMs aim to enable generalist locomotion control by supporting rapid adaptation via fine-tuning.
  • Multi-Layer Adaptive Frameworks (MLAFs) – modular, hierarchical architectures that span from HL planning to LL control, with each layer producing interpretable intermediate outputs. MLAFs are designed to integrate, substitute, and coordinate diverse policies, enabling flexible and adaptive responses across tasks and embodiments.
In the following subsections, we present detailed conceptual analyses of BFMs and MLAFs, evaluate their potential to realise a unified framework, and subsequently discuss fundamental challenges inherent to existing frameworks.
IV-A Bipedal Foundation Models
Inspired by Robot Foundation Models [60, 61], BFMs are large-scale, general-purpose models tailored for bipedal locomotion. Trained on diverse datasets spanning various morphologies, terrains, and objectives, BFMs enable reasoning about novel scenarios and adaptation with minimal task-specific supervision. A second-stage DRL fine-tuning process further adapts these models to specific platforms, supporting versatile locomotion behaviours such as walking and turning. Building upon the end-to-end reinforcement learning methods introduced in Section II, which typically focus on task-specific policy training, BFMs retain RL as a core adaptation mechanism, but shift toward a unified and generalisable paradigm by introducing shared motion priors across tasks and morphologies. This reuse of RL in a fine-tuning role allows BFMs to ground HL priors into executable LL policies, while improving sample efficiency and transferability.
The BFMs architecture includes a multi-modal embedding module, a latent backbone, and an action decoder. The embedding encodes morphology, task specs, and optional environment input into a shared space. The backbone converts these into a temporally structured latent representation conditioned on the task. The decoder outputs LL control commands adapted to the robot. Similar modular structures appear in RT-1 [62] and RT-2 [63], which map multimodal inputs to discrete actions using a shared transformer backbone.
The model first leverages IL to pre-train generalisable motion priors from diverse demonstrations, resulting in latent representations that facilitate scenario reasoning and transfer across different robot morphologies. DRL then fine-tunes these priors by aligning them with the physical constraints, action spaces, and dynamics of specific platforms. Rather than learning from scratch, DRL enables efficient adaptation to new tasks. Lightweight strategies such as ControlNet-style modular layers such as AdaptNet [64]—allow task-specific adjustments with minimal parameter updates.
Beyond modular adaptation, recent works have demonstrated that large-scale DRL fine-tuning significantly enhances generalisation and capability transfer. FLaRe [65] fine-tunes behaviour-cloned policies with sparse rewards to achieve high success rates on long-horizon mobile manipulation tasks, including generalisation to novel embodiments. Complementarily, MOTO [66] applies a model-based DRL framework that incorporates on-policy value expansion and epistemic uncertainty estimation, enabling effective offline-to-online adaptation from image observations.
Other frameworks further highlight this paradigm. AdA [67] demonstrates that scaling DRL with memory-based policies allows for in-context adaptation to novel embodied environments. RoboFUME [68] combines VLM-based reward classification with DRL fine-tuning to autonomously adapt pre-trained manipulation policies without environment resets. Collectively, these approaches underscore DRL not only as a fine-tuning tool, but as a central mechanism for grounding foundation model priors into executable, platform-specific control policies.
While DRL enables flexible and powerful adaptation of pre-trained motion priors, applying it in bipedal foundation models introduces several challenges. Fine-tuning often demands carefully shaped rewards, frequent environment resets, and large amounts of interaction data—factors that are especially costly and risky on physical bipedal robots [68]. Additionally, DRL policies may overfit to narrow dynamics or forget generalisable priors acquired during pre-training [65, 67]. These issues constrain the robustness, efficiency, and safety of DRL-based adaptation in real-world settings. Meanwhile, existing datasets such as AMASS [69] and Motion-X [70] remain human-centric and require substantial retargeting effort [25]. The inference latency of large models also limits real-time deployment on hardware-constrained platforms [63].
IV-B Multi-layered Adaptive Models
Unlike BFMs, which rely on large-scale pre-training and shared latent representations, MLAMs adopt a modular hierarchical approach, extending conventional hierarchical frameworks in Section III with explicitly adaptive layers. MLAMs decompose the locomotion pipeline into three distinct adaptive tiers: (1) HL reasoning for task decomposition and goal interpretation, (2) mid-level planning for behaviour generation, and (3) LL control for precise command execution. This modularity allows each layer to be independently optimised, upgraded, or replaced without retraining the entire system. MLAMs distinguish themselves from conventional hierarchical approaches through their adaptive modularity, dynamic policy integration, and clear layer-wise semantic responsibilities.
A key distinguishing feature of MLAMs lies in its capacity to dynamically compose adaptive modules for each control tier. Instead of generalising through unified embeddings as in BFMs, MLAMs achieve scalability by dynamically assembling specialised policies. Each layer processes context-specific inputs and outputs interpretable commands. The HL reasoning layer typically uses symbolic reasoning or Large Language Models (LLMs) [71, 72] to parse commands into sub-tasks. For instance, Vision-Language Model Predictive Control [73] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning. By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution. The mid-level planner selects or synthesises motions via learned motion libraries [55, 41] or DRL-based planners [74].
The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands. These include locomotion primitives like walking and climbing [75], adaptive tracking controllers for whole-body tracking [56], and imitation-based skills such as kicking and dancing [22]. Controllers at this level utilise RL, IL, or model-based methods [9, 76] and operate at high frequency to ensure robust real-time execution across varied conditions. An example of MLAMs architecture is demonstrated by recent hierarchical systems using LLMs for quadrupedal tasks [75]. HL modules interpret instructions, mid-level planners select motion primitives, and LL controllers execute precise actions, enabling robust dynamic replanning and flexible real-world operation.
Despite these strengths, MLAMs deployment faces challenges distinct from BFMs. MLAMs depend on real-time multi-modal perception such as vision, IMU, force sensors, complicating data alignment across layers due to differing timescales and abstraction levels [76]. Additionally, HL reasoning modules like LLMs often incur computational latency from autoregressive inference, posing limitations for tasks needing rapid reactions [75]. Thus, integrating multi-modal perception and ensuring real-time execution are primary challenges for effective MLAMs implementation.
Taken together, MLAMs offer a complementary path to BFMs by emphasising modularity, interpretability, and adaptability through dynamic policy composition. While they avoid the scalability bottlenecks of monolithic models, their reliance on real-time multi-modal processing and multi-layer coordination introduces unique system-level challenges.
IV-C Fundamental challenges
The development of unified frameworks such as BFMs and MLAMs aim to address the challenges of bipedal locomotion in dynamic and unstructured environments. However, achieving such frameworks requires overcoming several fundamental obstacles. Among these, balancing generalisation and precision, as well as designing streamlined frameworks capable of handling complex tasks, represent two critical challenges. These challenges are explored in the following subsections.
IV-C1 Generalisation and precision
The challenge of simultaneously improving generalisation across diverse skills and precision in specific tasks, such as jumping precisely to a desired distance, remains a fundamental obstacle in realising unified frameworks like BFMs and MLAMs. For BFMs, which focuses on large-scale, versatile humanoid models, the difficulty lies in enabling the model to adapt to diverse tasks while maintaining precise control in dynamic environments. For MLAMs, which emphasizes task decomposition and hierarchical control, ensuring precision across sub-tasks while maintaining a HL general policy is equally challenging.
DRL has demonstrated potential in facilitating versatile locomotion skills [31]; however, challenges such as poor velocity tracking and issues with precise control often arise. While [53] shows that deep planning combined with model-based control can achieve precise velocity tracking, and [45] illustrates successful end-to-end control for precise jumping, the creation of a policy that effectively handles both diverse tasks and precise movements remains elusive. Furthermore, [51] introduces a foot constraint policy framework, enabling precise target tracking and accurate touchdown locations. Yet, there is still no framework that comprehensively addresses the dual demands of versatility and precision in locomotion.
For BFMs, these challenges are exacerbated by its reliance on diverse and large-scale datasets, which often lack the precision required for specific bipedal tasks. In contrast, MLAMs struggle to achieve precision across its hierarchical layers, where HL policies must generalise across tasks while LL controllers maintain stability and accuracy in dynamic scenarios.
The difficulty in simultaneously achieving precise control and a broad range of actions in bipedal locomotion using DRL stems from several factors:
  • Complex dynamics: Bipedal locomotion involves intricate dynamics, posing a significant challenge to maintaining both dynamic motion and precision.
  • Resource intensity: Executing diverse locomotion tasks requires considerable computational power and extensive data, necessitating high-quality hardware and efficient DRL algorithms.
  • Training conflicts: Training DRL systems to achieve both precision and versatility often leads to conflicts. Designing reward functions and training policies that satisfy both criteria is inherently complex.
These challenges underscore the need for innovative solutions that balance task generalisation with control precision, bridging the gap between BFM’s large-scale adaptability and MLAM’s hierarchical precision.
IV-C2 Streamlined frameworks capable of handling complex tasks
One of the critical challenges in unifying BFMs and MLAMs is designing frameworks that are both simple and capable of handling complex tasks. While streamlined frameworks reduce computational overhead and potential failure points, they may lack the flexibility to adapt to diverse and dynamic environments. On the other hand, complex frameworks with multiple components, such as foot planners or terrain-specific modules [35, 36], often introduce additional computational burden and coordination challenges.
For BFMs, the challenge lies in creating a model that leverages large-scale general representations without adding unnecessary layers of complexity. For MLAMs, the difficulty is in coordinating hierarchical decision-making across layers without introducing redundancy or inefficiency. Achieving a balance between simplicity and capability is essential for creating robust, unified frameworks.
The advantage of maintaining simplicity in the framework lies in its ability to streamline decision-making processes, thereby reducing computational overhead and potential points of failure. To achieve an optimal framework, advancements in several key areas are essential:
  • Robust and efficient DRL algorithms: Development of algorithms that can manage high-dimensional and continuous control problems more effectively.
  • Specialised neural network architectures: Design of neural architectures tailored for specific bipedal tasks, capable of processing extensive sensory data (e.g., visual and tactile inputs), similar to the innovations presented in [52].
  • Effective reward functions: Formulation of reward functions that more accurately guide the learning process towards achieving desired behaviours and strategic outcomes.
  • Advanced computational resources: Enhancement of computational capabilities to support more intensive training and faster inference, facilitating real-time decision-making in dynamic environments.
By focusing on these developmental areas, the potential to create a unified, efficient, and less complex framework for handling complex locomotion challenges in bipedal robots is significantly increased.
Figure 4: Towards a Unified Framework: the transition from the present frameworks (end-to-end and hierarchical frameworks) to a unified framework capable of handling diverse locomotion tasks. The present state is characterised by limitations such as fine-tuning requirements, limited gait patterns, restricted terrain walking, and blind locomotion. To bridge this gap, future developments must focus on generalisation, precision, streamlined frameworks, vision-based learning, and zero-shot sim-to-real capabilities. These advancements will lead to the creation of two potential unified models: (1) a Multi-Layered Adaptive Model grounded in the hierarchical framework and (2) a Bipedal Foundation Model built upon the end-to-end framework. The unified framework will ultimately support diverse locomotion tasks in complex environments.
V Future Directions
Section II and III introduced two mainstream learning-based frameworks, HFM and MLAM, and analysed their contributions toward achieving a unified framework.While Section IV provides a clearer understanding of the current state and the existing gap between these frameworks and the expected unified framework, further discussion is needed to outline specific directions for achieving this goal.
In this section, we outline the essential components and future research directions necessary to bridge this gap. Specifically, we identify key elements that will drive the evolution from current framework limitations to a fully integrated unified framework. A summary of these fundamental elements is provided in Fig. 4.
V-A Future directions for bipedal robots
V-A1 Loco-manipulation tasks
With the rapid development of bipedal robotics, modern bipedal robots are now equipped with upper-body structures to enhance their capabilities beyond basic locomotion tasks. These advancements enable robots to perform upper-body-assisted locomotion and loco-manipulation tasks, expanding their adaptability in complex environments.
Building on this evolution, upper-body-assisted locomotion enhances movement stability by integrating upper-lower body coordination, enabling robots to navigate challenging terrain more effectively. Beyond movement, manipulation tasks focus on precise upper-body control, requiring dexterous motion planning for object interaction and task execution. At the highest level of complexity, loco-manipulation integrates both locomotion and manipulation, demanding multi-contact interactions, dynamic balance maintenance, and real-time adaptation to unpredictable environments. These tasks necessitate full-body coordination and task-driven motion planning, making them essential benchmarks for advancing unified frameworks in humanoid robotics.
Tasks include fundamental locomotion movements like flat walking and stepping over obstacles, progressing to upper-body-assisted locomotion challenges such as climbing ladders, wall-assisted walking, jumping over obstacles, and grasping for support. Beyond locomotion, precise upper-body control is essential for manipulation tasks like turning valves, driving vehicles, using tools, and connecting fire hoses. The most complex scenarios demand full-body coordination for loco-manipulation tasks such as carrying objects while walking, opening doors while navigating, and executing disaster-response operations. These benchmarks are critical for developing robust and adaptable humanoid systems, addressing key challenges posed in DARPA’s robotics initiatives [39].
Therefore, loco-manipulation presents opportunities for humanoid robots to excel beyond purely bipedal capabilities. Few studies have addressed this integrated task; one such study [77] demonstrated a ’box transportation’ framework. This framework decomposes the task into five distinct policies, each addressing different aspects of the transportation process. However, this approach lacks efficiency and does not incorporate vision-based information, suggesting substantial room for improvement. Moreover, the challenges of managing mobile tools like scooters [78] or dynamically interacting with objects such as balls [79] introduce further complexities.
To address these challenges, a hierarchical control approach could provide a more structured and adaptable solution. Decomposing loco-manipulation into multiple layers would simplify task execution by allowing precise, modular control over individual components [53]. This method enables manual fine-tuning of task-specific controllers, ensuring better stability and adaptability to environmental variations. Additionally, a structured hierarchical approach facilitates coordinated whole-body responses, improving the efficiency of loco-manipulation interactions.
Alternatively, an end-to-end learning framework could offer a more generalised and scalable solution. By leveraging task randomisation and structured curriculum learning, bipedal robots could progressively learn loco-manipulation skills in a unified policy [48, 36, 31]. During training, such policies can also learn human-like movements from motion capture data [80, 27, 25], offering promising solutions for future integrated loco-manipulation tasks within a single, versatile policy.
V-A2 Bridging the gap from simulation to reality
Due to the large number of interactions needed for DRL algorithms, training directly on robots can lead to costly damage to hardware and the environment. Consequently, training a policy in the simulation which offer a safe and cost-effective environment for developing robotics policies, and then deploying it to the hardware illustrates significant potential and efficiency. Bridging this gap is a key step toward achieving a unified framework, as it allows for the implementation of more versatile and robust systems that can handle diverse tasks across different terrains. Furthermore, the ultimate goal of achieving zero-shot transfer, where a policy trained in simulation can be deployed without additional real-world fine-tuning, directly contributes to the unified framework’s adaptability and scalability. This capability would make the framework far more practical for handling the wide range of locomotion tasks expected from future bipedal robots.

### v5

@@ I Introduction
Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [14, 15, 16]. Early methods, such as the Linear Inverted Pendulum Model (LIPM) [17], allowed for simplified representations of the dynamics involved in bipedal movement, enabling easier analysis and control. As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion. Advanced methods such as Model Predictive Control (MPC) [18, 19] and Trajectory Optimisation (TO) [20, 21, 14] exploit a predefined dynamics model to solve constrained optimal-control problems planning footsteps, CoM motion, and contact forces. While model-based approaches offer rapid convergence and predictive capabilities, they often fall short in dynamically complex and uncertain environments, where adaptability is key.
@@
To provide a clear benchmark for answering this question, we first define a unified framework as a control architecture capable of supporting consistent learning, transfer, and execution across arbitrary robot morphologies, terrains, and task definitions. The ultimate aim is to develop genuinely generalist locomotion policies that achieve robust zero-shot generalisation to previously unseen scenarios, embodiments, and objectives.
With this definition as our guide, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical. These frameworks enable robots to generate diverse gaits, including walking, running, and turning, while also allowing them to navigate complex terrains, such as stairs, slopes, and deformable surfaces.
@@
Figure 2: Classification of DRL-based control schemes. The approaches are broadly categorised into two main paradigms: End-to-end frameworks, which learn a single policy from sensory inputs to motor commands, and Hierarchical frameworks, which decompose the control problem into multiple levels. Within the end-to-end paradigm, a key distinction is between Reference-free learning (learning from scratch) and Reference-based learning (tracking a predefined motion). Hierarchical structures include Hybrid control schemes, which synergistically combine learned components with traditional model-based controllers.
Current progress shows that we are still far from achieving a unified framework, despite two mainstream types of framework structure. Although bipedal locomotion tasks have evolved significantly in the past decade. Earlier frameworks primarily focused on basic walking stabilisation and simple terrain adaptation, whereas recent developments have introduced complex tasks such as dynamic stair climbing, parkour-inspired locomotion, and loco-manipulation. Notably, the DARPA Robotics Challenge (DRC) has established a structured benchmark for bipedal humanoid locomotion and manipulation [42], emphasizing real-world deployment and robustness.
@@
The survey is organised as follows: Section II focuses on end-to-end frameworks, categorised by learning approaches. Section III details hierarchical frameworks, classified into three main types. Section IV presents specific limitations and challenges, highlighting their connection to the previously surveyed content. Section V addresses potential future pathways, opportunities, and two proposed conceptual models that extend the end-to-end and hierarchical paradigms. Finally, Section VI concludes the survey.
@@ II End-to-end framework
TABLE I: Summary and comparison of reference-based and reference-free Learning approaches for end-to-end framework. The dashed line in the implementation flow chat refers to optional.00footnotetext: Forward Walk involves the bipeds walking straight ahead. Unidirectional Walk enables the bipeds to move both forward and backward within a range of desired velocities. Omni-Walk grants the bipeds the ability to walk in any direction. Versatile Walk allows the bipeds to walk forward, backward, turn, and move sideways, providing extensive movement capabilities. Periodic Motions entails the execution of various repeated gait patterns, such as walking, hopping, or galloping. Versatile Jump refers to jump towards different desired targets. Versatile Motions cover performing a broad array of motions, both periodic and aperiodic such as jumping.
Methods
Works
Capabilities
Characteristic
Advantages and Disadvantages
Implementation Flow Chart
Residual learning
[45]
[46]
Forward walk
undirectional walk
Add residual term to the known motor positions at the current time step.
A: Fast convergence speed D: Require high-quality predefined reference, limit to specific motions, and lack robustness to complicated terrains.
Guided learning
[47]
[44]
[48]
[35]
Forward walk
Versatile walk
Versatile jump
Versatile motions
Mimic the predefined reference and directly specifies joint-level command.
A: Accelerate the learning process and robust to the terrains. D: Limited to the predefined motions and lack adaptability to unforeseen changes in environment.
Reference-free learning
[32]
[49]
[33]
Periodic motions
Stepping stones walk
Visual walk
Learn locomotion skills from scratch without any prior knowledge.
A: High potential for gait exploration, high robust to the complicated terrain. D: Requires intensive reward shaping for gait patterns and relatively expensive computational training resources.
@@ III Hierarchy framework
Control Scheme
Works
Module
Characteristic
Advantages and Disadvantages
Deep Planning Hybrid Scheme
[36]
[56]
[57]
Deep planning + ID
Deep planning + ID-QP
Deep planning + WPG
HL policy is learned to guide the LL controller to complete locomotion and navigation tasks.
A: Enhanced command tracking capabilities, generalised across different platforms, sampling efficiency, and robust. D: Complicate system and communication between layers, require precise model, lack generalisation regarding different tasks.
Feedback DRL Control Hybrid Scheme
[58]
[38]
[39]
Gait library + feedback policy
Footstep planner + feedback policy
Model-based planner + feedback policy
LL feedback policy receives non-learned HL planner as input to achieve locomotion skills.
A: Short inference times, robust, navigation locomotion capabilities, interpretability. D: Complicated system and communication between layers, reducing sampling efficiency.
Learned Hierarchy Framework
[25]
[41]
[59]
HL policy + LL policy
HL policy + LL policy
HL policy + LL policy
Both HL planner and LL feedback controller are learned. LL policy focuses on basic locomotion skills; on the other side, HL policy learn navigation skills.
A: provides layer flexibility, where each layer can be independently retrained and reused; alleviates the challenges associated with training an end-to-end policy. D: inefficienty sim-to-real, complicated interface between layers, training expensively.
@@ III-A Deep planning hybrid scheme
Figure 3: Hierarchy Control Scheme Diagram: presents a hierarchical control framework for a bipedal robot, consisting of a basic scheme and three variations. (1) Basic Scheme: The framework starts with a Task Command, followed by a HL Planner, and a LL Controller, which ultimately drives the Robot. Each module can be replaced by a learned policy, introducing adaptability in different control layers. (2) Variations (from left to right): (a) A deep planning hybrid scheme, where the HL planner is learned. (b) A Feedback DRL control hybrid scheme, with a learned $L$ controller. (c) A Learned hierarchy control scheme, where both layers are learned.
@@ III-C Learned hierarchy framework
IV Limitations and Challenges
The end-to-end and hierarchical frameworks detailed in Sections II and  III represent the state-of-the-art in DRL-based bipedal locomotion, demonstrating significant capabilities on specific tasks. However, a significant gap persists between these specialised successes and the realisation of a truly unified framework as defined in Section  I.
Evaluating existing approaches against this definition reveals that neither the end-to-end nor hierarchical control paradigms currently meet these criteria. The path to a unified framework is hindered by several interconnected issues. At a foundational level, a primary difficulty involves the limitations and challenges in achieving both generalisation and precision (Section IV-A). This is further complicated by the practical barrier of the ‘reality gap’ in transferring policies from simulation to physical robots (Section IV-B). Ultimately, these issues culminate in the critical challenges of ensuring safety and interpretability for robust deployment in real-world, safety-critical situations (Section IV-C).
IV-A Generalisation and precision
A central challenge in applying DRL to bipedal locomotion is the need to simultaneously achieve high generalisation across diverse skills and transverse to all kinds of terrains, and high precision in specific tasks. This remains a fundamental obstacle to realising truly unified and capable frameworks.
This capability gap is evident in the current literature. Many approaches excel at generalisation, demonstrating policies that enable versatile skills like walking and jumping [44, 35] and can transfer to different terrains [49, 33, 63]. However, these generalized policies often lack the fidelity required for high-precision tasks, such as exact foot placement [38, 33, 54] or maintaining a specific velocity with minimal error [52]. Conversely, controllers specialized for narrow domains can achieve exceptional precision, as seen in jumping to a precise target [48], yet they cannot generalize these capabilities to a broader range of tasks. Thus, the development of a single, unified framework that concurrently exhibits both broad competency and high fidelity remains a largely unresolved challenge.
This difficulty in uniting generalisation and precision is not arbitrary but stems from several key limitations inherent in current DRL paradigms, whether related to framework design, task formulation, or the training process itself:
  • Limited terrain and gait patterns: The failure to generalise is often a direct result of training on insufficiently diverse environments or with a restricted set of behaviours. Models trained on limited terrain are brittle when faced with novel surfaces, while a limited gait pattern library prevents them from adapting to tasks requiring new motor skills.
  • Poor command tracking: The learning signals for generalisation and precision are often in direct conflict. generalisation requires permissive signals that allow the robot to adapt to varied terrains or recover from perturbations. In contrast, precision demands restrictive signals that minimize command tracking error. Faced with these opposing objectives, a single policy is forced to compromise, which often leads to poor command tracking and the sacrifice of adaptability in favour of rigid, high-fidelity execution [34, 35].
  • Inefficient sampling: Underpinning the difficulty of solving both problems simultaneously is the inefficient sampling of most DRL algorithms [40, 64, 65]. This problem is severely exacerbated in tasks that depend on sparse rewards, where feedback is infrequent and often only provides support for the success of the final task, thus the task, hence the community’s frequent resort to reward shaping and curriculum learning [53, 32, 58, 66]. Consequently, the immense amount of data required for an agent to explore, stumble upon a successful strategy, and then refine it for both a diverse skill set for generalisation and the fine-grained control needed for precision is often computationally prohibitive, motivating massive parallel simulation just to make training tractable [67, 68, 25, 26].
  • High-quality data scarcity: As highlighted in our discussion in Section LABEL:subsec:guided_learning, the scarcity of high-quality demonstrations is a key bottleneck.Such data provides essential guidance for DRL, enabling policies to learn physically feasible and natural-looking gaits while bypassing extensive, unsafe exploration [26, 25, 69]. This scarcity stems from the fact that scalable human data is difficult to transfer due to embodiment mismatch [70, 71, 59], while generating feasible synthetic data via trajectory optimisation is often computationally expensive [20, 21].
These fundamental limitations give rise to common algorithmic challenges, such as the need for complex reward engineering, and are directly embodied in the designs of the field’s dominant control architectures. End-to-end frameworks attempt a holistic solution, learning a single, monolithic policy that must implicitly resolve all challenges simultaneously. While this approach can yield highly versatile and dynamic behaviours [63], it directly confronts the immense difficulty of exploration from sparse rewards and the struggle of reconciling conflicting training objectives within the system’s unstable dynamics. This often results in a notable lack of the fidelity and precision that hierarchical systems can enforce [35]. Conversely, hierarchical frameworks, for example, are a direct architectural response to the lack of skill compositionality. By employing a “divide and conquer” strategy, they use a HL policy to sequence a library of LL, often model-based, controllers. This structure is designed to enforce precision and manage the complex dynamics at a lower level [21, 54, 39]. However, as noted, this results in a brittle system, imposing a strong prior that constrains the policy’s freedom and limits its ability to generalise to situations not anticipated by the handcrafted controller [38].
IV-B Reality gap: From simulation to the physical world
Another challenge hindering the deployment of DRL policies on bipedal robots is the simulation-to-reality (sim-to-real) gap. This refers to the significant discrepancy between a policy’s performance in a physics simulator and its performance on actual hardware. This gap is a critical obstacle because training directly on physical robots is often impractical. The millions of environmental interactions required for DRL would lead to accelerated mechanical wear, risk of catastrophic failure, and require constant human supervision. While simulation offers a safe and efficient alternative, the ultimate goal of “zero-shot” transfer, where a policy works perfectly without any real-world fine-tuning, is rarely achieved.
A large body of research validates impressive locomotion skills purely within simulation, without attempting transfer to a physical system [25, 26, 72, 73, 34]. Even when transfer is successful, it often comes with compromises. Many successful transfers are not truly “zero-shot” and rely on a subsequent phase of extensive real-world fine-tuning or manual parameter tuning [74, 46]. In cases where policies do transfer without fine-tuning, they often exhibit a noticeable degradation in performance, where the robustness and agility seen in simulation are significantly lower in the real world [49, 33, 75].
This gap is caused by unavoidable differences between the virtual and physical worlds, which are especially problematic for dynamically unstable bipedal robots.
  • Robot dynamics modelling and actuation: Simulators struggle to perfectly replicate the complex dynamics of a physical bipedal robot, whose inherent instability makes it particularly sensitive to modeling errors. Factors such as motor friction, gear backlash, and precise link inertia are often simplified.
  • Contact and terrain modelling: Accurately simulating the intermittent foot-ground contact that bipeds rely on is extremely difficult. A mismatch between simulated and real-world friction or surface properties can cause unexpected slips or bounces, leading to a loss of balance.
  • Sensing and state estimation: A simulated robot has access to perfect, noise-free state information. In the real world, these states must be estimated from noisy sensors like IMUs and joint encoders [46, 74]. For a bipedal robot, precise state estimation is critical for maintaining balance.